​(New project analysis​) Garcia's Truckin' Inc. is considering the purchase of a new production machine for $150,000. The purchase of this machine will result in an increase in earnings before interest and taxes of ​$40,000 per year. To operate the machine​ properly, workers would have to go through a brief training session that would cost ​$5,000 after taxes. It would cost​$6,000 to install the machine properly.​ Also, because this machine is extremely​ efficient, its purchase would necessitate an increase in inventory of $25,000. This machine has an expected life of 10 ​years, after which it will have no salvage value.​ Finally, to purchase the new​ machine, it appears that the firm would have to borrow​ $100,000 at 9 percent interest from its local​ bank, resulting in additional interest payments of ​$9,000 per year. Assume simplified​ straight-line depreciation and that the machine is being depreciated down to​ zero, a 36 percent marginal tax​ rate, and a required rate of return of 11 percent.

a. What is the initial outlay associated with this​ project?

b. What are the annual​ after-tax cash flows associated with this project for years 1 through​ 9?

c. What is the terminal cash flow in year 10

​(what is the annual​ after-tax cash flow in year 10

plus any additional cash flows associated with the termination of the​ project)?

d. Should the machine be​ purchased?

a. What is the initial outlay associated with this​ project?

​$_ ​(Round to the nearest​ dollar.)
Answer #1

a)

initial outlay:

cost of the machine = 150000

installation cost = 6000

Training cost    = 5000

increase in working capital = 25000

Initial outlay = 186,000

b)

depreciation = 156,000 / 10 = 15,600 (cost + installation expenses = 156000)

annual after tax cash flows

EBIT 40000
less : interest (9000)
less: Depreciation (15,600)
net income 15400
less : Taxes @36% (5544)
earnings after tax 9856
Add: Depreciation 15600
Free cash flows 25456

c)

Terminal cash flow:

in last year i.e.,year 10 in addition to free cash flow we recover initial working capital investment

so terminal cash flow = 25456 + 25000 = $50,456

d)

we have to calculate NPV of the project

NPV = present value of free cash flows - initial cash outflow

present value of free cash flows:

PVIFA ( r = 11% ; n = 9) = 5.537047

PV factor ( r = 11% ; n = 10) = 0.352184478

present value = (25456 x 5.537047) + (50456 x 0.352184478) = 158720.90

NPV = 158720.90 - 186000 = -27279.10

Since NPV is negative we should not purchase the machine.